Problem 10.16

You are graduating in two years. You want to invest your current savings of $6,100 in bonds and use the proceeds to purchase a new car when you graduate and start to work. You can invest the money in either bond A, a two-year bond with a 3.23 percent annual interest rate, or bond B, an inflation-indexed two-year bond paying 1.03 percent real interest above the inflation rate (assume this bond makes annual interest payments). The inflation rate over the next two years is expected to be 1.46 percent. Assume that both bonds are default free and have the same market price. Which bond should you invest in? (Round answers to 2 decimal places, e.g. 15.25%)

Nominal interest rate for Bond A

%
Nominal interest rate for Bond B

%
Invest in

Bond B or Bond A

Problem 10.18

Healthy Potions, Inc., is considering investing in a new production line for eye drops. Other than investing in the equipment, the company needs to increase its cash and cash equivalents by $10,000, increase the level of inventory by $46,000, increase accounts receivable by $25,000, and increase accounts payable by $5,000 at the beginning of the project. Healthy Potions will recover these changes in working capital at the end of the project 9 years later. Assume the appropriate discount rate is 15.0 percent. What are the present values of the relevant investment cash flows? (Round answer to 2 decimal places, e.g. 15.25.)

Present value $
Answer #1

Answer:

Nominal interest rate for Bond A 3.23
Nominal interest rate for Bond B 2.49
Invest in Bond -A

Nominal interest rate for Bond A = 3.23 %.....(usually stated interest rate is the nominal interest)

Nominal interest for Bond B = real interest rate + inflation rate => 1.03% + 1.46% =>2.49%...(formula method can be used if real rate and inflation rate are given).

Investment shall be made in Bond A.....(since it offers a higher interest, given that they have same market price and risk).

Present Value $21,603.18

Cash outflow at the beginning of the project = Increase in cash and cash equivalents + increase in the level of inventory + increase in accounts receivable - increase accounts payable
= 10000 + 46000 + 25000 - 5000
= $76,000

As Healthy Potions will recover these changes in working capital at the end of the project 9 years later.

So, Present Value of Cash inflow = 76,000 / (1+r)9
= 76000 / (1.15)9
= 76000/3.518
= $21,603.18